# Towards improved property prediction of two-dimensional (2D) materials using many-body Quantum Monte Carlo methods

### v1

@@ III.3.1 Bilayer Phosphorene
Layered phosphorene allotropes have garnered a great deal of interest because of its layer dependent properties exemplified with band gaps of black phosphorenes varying from 2.26 at the monolayer limit to 0.3 eV for the bulk [260, 261, 262]. This is understood to be driven by hybridization of valance pz orbitals between the adjacent layers, which leads interlayer interactions of layered phosphorenes to be not purely dispersive as compared to those of bilayer graphene interpreted as a typical vdW system.
@@
The distinct interlayer interaction of phosphorenes from the work of Shulenburger et al. [259] is revealed through Fig. 13 presenting DMC interlayer binding energy curves for AA- and AB-stacked bilayer black phosphorenes. There is a significant energy difference of interlayer interactions ( 18 meV per atom) and interlayer separations ( 0.6 Å) for AA- and AB-stacked bilayer black phosphorenes, not consistent with the vdW interaction between two homogeneous slabs. This is contrast to the case of bilayer graphene showing difference of 6 meV/atom and 0.1 Å for interlayer binding energies and separations, respectively [256, 257]. Furthermore, the charge density redistributions computed with DMC for the two stacking modes shows a delicate difference in the interlayer region where AB and AA yield the charge depletion and accumulation, respectively. This feature demonstrates the nature of interlayer interactions in black phosphorenes to be complicated, not characterized by purely dispersive interactions.
@@
Additionally, blue phosphorene, a 2D phosphorus allotrope with a puckered honeycomb structure, shows interlayer binding properties analogous to those of black phosphorene, according to our DMC study reported in Ref. 264. Furthermore, as the interlayer distance decreases, a blue phosphorene bilayer could exhibit a semiconductor-to-metal transition with the metallic state being manifested at short interlayer distances comparable to the intralayer bond length [263], which is observed for the A1B-1 stacking mode, one of five possible bilayer stacking modes of a puckered honeycomb structure [265]. A successful capture of this feature requires a well-balanced description of long-range dispersive forces and short-range chemical bondings with DMC because DFT calculations based on different exchange-correlation functionals show large inconsistency in predicting equilibrium interlayer binding energies and interlayer separations (as displayed in Fig. 14.)
@@
Figure 14(b) presents the DMC interlayer binding curve as a function of the interlayer distance, along with the corresponding DFT ones, for which the lattice constant optimized at a given interlayer distance was used (see Fig. 14(a)). The DMC binding curve is seen to possess two minima with their energy difference being 76(1) meV per atom. Among the DFT functionals considered here, the SCAN+rVV10 (SCAN plus vdW correction) [80] is found to produce the results in the best agreement with DMC in terms of the two-minima feature and the relative energetics between them. The local minimum at a short interlayer distance and the global minimum at a longer distance turn out to have a metallic and a semiconducting band structure, respectively. The metastable metallic minimum at a short interlayer distance is understood to arise from the interlayer hybridization between pz orbitals of the highest occupied valance bands of the two monolayers as well as intralayer charge redistribution. Random phase approximation (RPA) calculations of Arcudia et al. [265] also revealed two energy minima in the interlayer binding energy curve with different electronic phases for an A1B-1-stacked blue phosphorene bilayer. Contrary to the DMC results, however, the RPA study predicted that the metallic minimum at a short distance was energetically favored over the semiconducting one at a longer distance. This discrepancy is understood to be due to a well-known RPA limitation in describing short-range interactions which play crucial roles in stabilizing the metallic phase. Additional information regarding this work can be found in Ref. 264.
@@ III.3.2 Bilayer Arsenene
Similar to phosphorene, arsenene (As) is another monoelemental 2D material that displays promising electronic properties[267, 268, 269, 270]. Monolayer arsenene can exist in the washboard (w-As) honeycomb structure (similar to black phosphorene) and the buckled (b-As) honeycomb structure (similar to blue phosphorene). For bilayer arsenene, the ordering of energetic stability and the electronic properties have not been clearly established and various local, semilocal, vdW corrected, and hybrid functionals have been used to study bilayer arsenene [271, 27, 272, 273, 274]. Due to the fact that this system is weakly bonded, vdW functionals are a viable method to describe the complex physics of bilayer arsenene. In order to benchmark these vdW functionals and obtain an accurate benchmark of the stability order of possible stacking configurations of bilayer arsenene, we performed DMC calculations [266].
In this work, we performed DMC and DFT (using local, semilocal and vdW density functionals and functionals with semi-empirical vdW corrections) calculations of the binding energy and interlayer distance of w-As and b-As in AA- and AB- stacking configurations (see example in Fig. 15). Our DMC results revealed the AA-stacking to be lower in energy than the AB-stacking for b-As. We also find that the layer-layer interactions are dispersive since the energy changes by 22 and the interlayer distance changes by 0.1 Å when going from AA- to AB-stacking. We find that for b-As, the interlayer distance changes by 0.65 Å when going from the AB- to AA- stacking, which most likely signifies a complex layer-layer interaction. When benchmarking our DMC results along with DFT, we find that vdW density functionals (i.e., DF [78]) can reproduce DMC energetics, but structural parameters are better described by semi-empirically corrected functionals such as the D3-G method of Grimme [83]. Additional information regarding this work can be found in Ref. 266.
@@ III.3.3 Bilayer Graphene and Graphyne
Low-dimensional carbon allotropes, -graphyne was predicted to possess weaker DFT binding energy for the bilayer structure than bilayer graphene, but has attracted a great deal of attention due to its larger honeycombs structure than graphene [257]. However, DFT studies of the bindings of graphynes failed to confirm the most stable stacking mode because of the strong dependence of vdW-corrected XC functional on computed DFT binding energy. Those uncertainties on choice of XC functional for DFT and the weak binding energy of graphynes leads to conclude that a more precise method is imperative for accurate prediction of stable stacking mode of graphynes and their binding energies.
Figure 16 shows our DMC binding energies of AB and Ab stacking mode of bilayer -graphyne as a function of the layer spacing. Negligible difference in equilibrium binding energy between AB and Ab mode from DMC indicates the difficulty of synthesizing a pristine AB or Ab stacking -graphyne structure. It is interesting to see both AB (23.2(2) meV per atom) and Ab modes (22.3(3) meV per atom) of graphyne show larger equilibrium binding energies than AB stacking bilayer graphene (17.8(3) meV per atom), which can be understood by the existence of other interlayer binding nature than the weak vdW interaction in a sp-sp2 hybridized graphyne form. Due to this difference of interlayer binding nature between sp2-bonded and sp-sp2 hybridized carbon networks, it is found that vdW-corrected DFT functionals including DFT-D2 [82], vdW-DF, and rVV10 significantly underestimate the binding energy of bilayer -graphynes while an overestimation of the binding energy was shown on the pristine sp2-bonded graphene. Among the vdW-corrected DFT functionals, binding energies computed by rVV10 functional show the closest result to corresponding DMC result, while vdW-DF provides the closest charge density distributions. This inconsistent trend indicates the importance of an accurate description for both of dispersion and density correction in vdW corrected DFT functionals, which can provide guidelines to improve vdW descriptions for the future Kohn-Sham scheme. More information can be found in Ref. 257.
In addition to our work presented in this section, Mostaani et al. [256] computed the DMC binding energy of bilayer graphene to be 11.5(9) meV/atom and 17.7(9) meV/atom for AA and AB-stacking respectively and Krongchon et al. [275] utilized DMC and tight binding models to accuractely describe the registry-dependent potential energy and lattice corrugation of twisted bilayer graphene, a highly correlated system.
@@ III.3.4 Bilayer and Bulk TiS2
The family of vdW density functionals has been largely successful for a wide variety of systems, including 2D materials. Due to the success of this approach, there have been numerous improvements to the original vdW-DF functionals, where an emphasis has been placed on incorporating semilocal exchange [277]. Despite these improvements, highly accurate benchmarking calculations, such as those performed with DMC, are crucial in paving the way to developing a general purpose vdW density functional that can accurately reproduce fundamental properties such as the total energy and the electron density. In this specific work, we chose to benchmark the properties of bulk and bilayer TiS2 with a variety of local, semilocal, and vdW DFT functionals against DMC [276]. TiS2 (part of the dichalcogenide family) is a system of interest due to its potential applications as a cathode material for Li-ion batteries [278] and because it lies within a phase transition from a narrow gap semiconductor to a semimetal, adding complexity and novelty to the noncovalent interactions [279].
After performing these benchmarking calculations (depicted in Fig. 17), we found a close relationship between the accuracy of the interlayer distance and binding energy. We find that more recently developed vdW functionals such as vdW-DF-optB88 perform well for both properties. In terms of the response of the electron density to binding, we find that functionals such as LDA and PBEsol [280] outperform the vdW functionals and can correctly describe the interlayer charge accumulation. vdW functionals such as vdW-DF-C09 [277], which was developed on purely theoretical grounds, performs the best in terms of simultaneously reproducing the DMC energy and electron density. This highlights the need for a theory-driven path forward to develop a fully predictive and consistent vdW functional, with highly accurate benchmarks from methods such as DMC paving the way. Additional information can be found in Ref. 276.
@@ III.3.5 Bulk CrI3
In this section, we revisit the CrI3 system (discussed extensively in Section III.1.2) from a different perspective. As previously mentioned, CrI3 exhibits long-range magnetic ordering from few-layer to monolayer [16]. Bulk and few-layer CrI3 possess long-range interlayer and short-range intralayer interactions. The long-range vdW interlayer (noncovalent) interactions, which are due to strongly correlated electrons occupying d orbitals of Cr, and the competing intralayer correlations make this an extremely challenging system to model with approaches such as local or semilocal DFT. Certain vdW corrections in DFT can improve the accuracy of these interlayer forces, but the results can vary drastically depending on which vdW correction is employed.
In this work, we studied the binding properties of bulk CrI3 with DMC, with the goal of simultaneously describing the short- and long-range correlations and overcoming the shortcomings of DFT [281]. For the monoclinic bulk CrI3 structure, we calculated the interlayer separation distance to be 6.749(73) Å with DMC, which is in excellent agreement with the experimental value of 6.623 Å [282]. We also estimated the interlayer binding energy to be between 14.3 to 17.9 meV/Å2. We benchmarked several DFT functionals with and without vdW corrections against our DMC results for interlayer separation distance and binding energy (See Fig. 18) and we found that vdW-DF-optB88 [79] and vdW-DF-optB86b [283] are closest at reproducing our DMC values. In addition, we studied the bulk rhombohedral structure of CrI3 and found that the rhombohedral and monoclinic are within thermal energy differences of each other, which is in agreement with experiment [282]. We believe our DMC benchmark calculations can be useful for testing other promising current and future vdW corrected functionals. Further information of this work can be found in Ref 281.
@@ III.3.6 Graphene-supported Pt layers
Recent experimental realization of 2D layered structures of Pt atoms above a graphene surface [285, 286, 287] has opened an interesting avenue for exploration of their various morphologies, as well as the bonding nature of Pt-graphene complexes. In the sense that the layer formation is a consequence of a competition between non-directional metallic bondings among Pt atoms and their covalent and vdW bondings to graphene, DMC benchmarking is critical for accurate theoretical investigation of these layered systems with large degrees of freedom (including different Pt/C atomic ratios). As far as their geometries and relative energetics between different layered structures of Pt atoms are concerned, DFT results based on the SCAN+rVV10 functional are found to best agree with the corresponding DMC results.
Our DMC-benchmarked DFT calculations show that a (111)-packing structure, where each of the Pt atoms is adsorbed at a hollow site of graphene (see Fig. 19(c)), is more stable than the (100)-packing ones when forming the commensurate Pt-graphene complex at a Pt/C atomic ratio of 1/2. This can be understood by a significant lattice mismatch (>10 ) between pristine graphene and a freestanding (100)-packing Pt layer. Energetic stability of incommensurate Pt layers with different Pt/C atomic ratios were also investigated. Figure 19 displays DFT-SCAN+rVV10 formation energies of both flat and buckled incommensurate (100)-packing Pt monolayers (see Fig. 19(a)) as functions of the Pt/C atomic ratio. The lowest energy configuration is found at the ratio of 0.457 and its formation energy is lower than the corresponding energies for buckled (100)-packing monolayer with the atomic ratio of 1/2, which is in line with recent experimental findings [287]. However, the formation energy of this optimal incommensurate structure is still higher than that of the (111)-packing commensurate structure. Through its systematic investigation for various possible morphologies of layered Pt-on-graphene systems, this DMC-benchmarked study contributes to the expansion of a family of available metallic Pt layers. Additional information can be found in Ref. 284.

### v2

@@ III.3.1 Bilayer Phosphorene
Layered phosphorene allotropes have garnered a great deal of interest because of their layer-dependent properties exemplified with band gaps of black phosphorenes varying from 2.26 eV at the monolayer limit to 0.3 eV for the bulk.[310, 311, 312] This is understood to be driven by hybridization of valence pz orbitals between the adjacent layers, which causes interlayer interactions of layered phosphorenes to be not purely dispersive compared to those of bilayer graphene interpreted as a typical vdW system.
@@
The distinct interlayer interaction of phosphorenes from the work of Shulenburger et al. [313] is shown in Fig. 13, presenting DMC interlayer binding energy curves for AA- and AB-stacked bilayer black phosphorenes. There is a significant energy difference of interlayer interactions (18 meV per atom) and interlayer separations (0.6 Å) for AA- and AB-stacked bilayer black phosphorenes, not consistent with the vdW interaction between two homogeneous slabs. This is in contrast to the case of bilayer graphene showing a difference of 6 meV/atom and 0.1 Å for interlayer binding energies and separations, respectively.[307, 308] Furthermore, the charge density redistributions computed with DMC for the two stacking modes show a delicate difference in the interlayer region where AB and AA yield the charge depletion and accumulation, respectively. This feature demonstrates that the nature of interlayer interactions in black phosphorenes is complicated, not characterized by purely dispersive interactions.
@@
Additionally, blue phosphorene, a 2D phosphorus allotrope with a puckered honeycomb structure, shows interlayer binding properties analogous to those of black phosphorene, according to our DMC study reported in Ref. 314. Furthermore, as the interlayer distance decreases, a blue phosphorene bilayer could exhibit a semiconductor-to-metal transition with the metallic state being manifested at short interlayer distances comparable to the intralayer bond length,[315] which is observed for the A1B-1 stacking mode, one of five possible bilayer stacking modes of a puckered honeycomb structure.[316] A successful capture of this feature requires a well-balanced description of long-range dispersive forces and short-range chemical bondings with DMC because DFT calculations based on different XC functionals show large inconsistency in predicting equilibrium interlayer binding energies and interlayer separations (as displayed in Fig. 14.) The balanced description of diverse interactions is a key feature achieved through DMC, which has been proven to exhibit the accuracy of CC calculations with single, double, and perturbative triple excitations (CCSD(T)), the gold standard in quantum chemistry, in systems such as the benzene dimer[317, 318] and methane-water.[319, 320] This often positions DMC as a reference method for relatively large molecular systems where the application of CCSD(T) is limited.[321, 322, 323, 324, 325]
@@
Fig. 14(b) presents the DMC interlayer binding curve as a function of the interlayer distance, along with the corresponding DFT ones, for which the lattice constant optimized at a given interlayer distance was used (see Fig. 14(a)). The DMC binding curve is seen to possess two minima with their energy difference being 76(1) meV per atom. Among the DFT functionals considered here, the SCAN+rVV10 (SCAN plus vdW correction) [80] is found to produce the results in the best agreement with DMC in terms of the two-minima feature and the relative energetics between them. The local minimum at a short interlayer distance and the global minimum at a longer distance turn out to have a metallic and a semiconducting band structure, respectively. The metastable metallic minimum at a short interlayer distance is understood to arise from the interlayer hybridization between pz orbitals of the highest occupied valence bands of the two monolayers as well as intralayer charge redistribution. Random phase approximation (RPA) calculations of Arcudia et al.[316] also revealed two energy minima in the interlayer binding energy curve with different electronic phases for an A1B-1-stacked blue phosphorene bilayer. Contrary to the DMC results, however, the RPA study predicted that the metallic minimum at a short distance was energetically favored over the semiconducting one at a longer distance. This discrepancy is understood to be due to a well-known RPA limitation in describing short-range interactions which play crucial roles in stabilizing the metallic phase. Additional information regarding this work can be found in Ref. 314.
@@ III.3.2 Bilayer Arsenene
Similar to phosphorene, arsenene (As) is another monoelemental 2D material that displays promising electronic properties.[326, 327, 328, 329] Monolayer arsenene can exist in the washboard (w-As) honeycomb structure (similar to black phosphorene) and the buckled (b-As) honeycomb structure (similar to blue phosphorene). For bilayer arsenene, the ordering of energetic stability and the electronic properties have not been clearly established, and various local, semilocal, vdW-corrected, and hybrid functionals have been used to study bilayer arsenene.[330, 27, 331, 332, 333] Because this system is weakly bonded, vdW functionals are a viable method to describe the complex physics of bilayer arsenene. To benchmark these vdW functionals and obtain an accurate benchmark of the stability order of possible stacking configurations of bilayer arsenene, we performed DMC calculations.[334]
In this work, we performed DMC and DFT (using local, semilocal, and vdW density functionals and functionals with semi-empirical vdW corrections) calculations of the binding energy and interlayer distance of w-As and b-As in AA- and AB-stacking configurations (see Fig. 15). Our DMC results revealed the AA stacking to be lower in energy than the AB stacking for b-As. We also find that the layer-layer interactions are dispersive because the energy changes by 22 and the interlayer distance changes by 0.1 Å when going from AA to AB stacking. We find that for b-As, the interlayer distance changes by 0.65 Å when going from the AB to AA stacking, which most likely signifies a complex layer-layer interaction. When benchmarking our DMC results along with DFT, we find that vdW density functionals (i.e., DF [78]) can reproduce DMC energetics, but structural parameters are better described by semi-empirically corrected functionals such as the D3-G method of Grimme.[83] Additional information regarding this work can be found in Ref. 334.
@@ III.3.3 Bilayer Graphene and Graphyne
Regarding low-dimensional carbon allotropes, -graphyne was predicted to possess weaker DFT binding energy for the bilayer structure than bilayer graphene but has attracted a great deal of attention due to its larger honeycomb structure than that of graphene.[308] However, DFT studies of the bindings of graphynes failed to confirm the most stable stacking mode because of the strong dependence of vdW-corrected XC functional on computed DFT binding energy. Those uncertainties in the choice of the XC functional for DFT and the weak binding energy of graphynes revealed that a more precise method is imperative for accurately predicting the stable stacking mode of graphynes and their binding energies.
Fig. 16 shows our DMC binding energies of the AB- and Ab-stacking modes of bilayer -graphyne as a function of the layer spacing. The negligible difference in equilibrium binding energy between AB and Ab modes from DMC indicates the difficulty of synthesizing a pristine AB- or Ab-stacking -graphyne structure. Both AB (23.2(2) meV per atom) and Ab modes (22.3(3) meV per atom) of graphyne show larger equilibrium binding energies than AB-stacked bilayer graphene (17.8(3) meV per atom). This can be attributed to the different nature of interlayer interactions in the sp-sp2 hybridized graphyne structure, as opposed to graphene whose interlayer coupling is primarily governed by weak vdW interactions. Due to this difference of interlayer binding nature between sp2-bonded and sp-sp2 hybridized carbon networks, it is found that vdW-corrected DFT functionals including DFT-D2, [82] vdW-DF, and rVV10 significantly underestimate the binding energy of bilayer -graphynes, while an overestimation of the binding energy was shown on the pristine sp2-bonded graphene. Among the vdW-corrected DFT functionals, binding energies computed by the rVV10 functional show the closest result to the corresponding DMC result, while vdW-DF provides the closest charge density distributions. This inconsistent trend indicates the importance of an accurate description for dispersion and density correction in vdW-corrected DFT functionals, which can provide guidelines to improve vdW descriptions for the future Kohn-Sham scheme. More information can be found in Ref. 308.
In addition to our work presented in this section, Mostaani et al. [307] computed the DMC binding energy of bilayer graphene to be 11.5(9) meV/atom and 17.7(9) meV/atom for AA and AB stacking, respectively, and Krongchon et al. [335] utilized DMC and tight binding models to accurately describe the registry-dependent potential energy and lattice corrugation of twisted bilayer graphene, a highly correlated system.
@@ III.3.4 Bilayer and Bulk TiS2
The family of vdW density functionals has been largely successful for a wide variety of systems, including 2D materials. Due to the success of this approach, there have been numerous improvements to the original vdW-DF functionals, where an emphasis has been placed on incorporating semilocal exchange.[336] Despite these improvements, highly accurate benchmarking calculations, such as those performed with DMC, are crucial in paving the way to developing a general-purpose vdW density functional that can accurately reproduce fundamental properties such as the total energy and the electron density. In this specific work, we chose to benchmark the properties of bulk and bilayer TiS2 with a variety of local, semilocal, and vdW DFT functionals against DMC.[337] TiS2 (part of the dichalcogenide family) is a system of interest due to its potential applications as a cathode material for Li-ion batteries [338] and has been debated to be either a heavily self-doped narrow gap semiconductor or a semimetal,[339] adding complexity and novelty to the noncovalent interactions.[340]
After performing these benchmarking calculations (depicted in Fig. 17), we found a close relationship between the accuracy of the interlayer distance and binding energy. We find that more recently developed vdW functionals such as vdW-DF-optB88 [79] perform well for both properties. In terms of the response of the electron density to binding, we find that functionals such as LDA and PBEsol [341] outperform the vdW functionals and can correctly describe the interlayer charge accumulation. The vdW functionals such as vdW-DF-C09,[336] which was developed on purely theoretical grounds, perform the best in terms of simultaneously reproducing the DMC energy and electron density. This highlights the need for a theory-driven path forward to develop a fully predictive and consistent vdW functional, with highly accurate benchmarks from methods such as DMC paving the way. Additional information can be found in Ref. 337.
@@ III.3.5 Bulk CrI3
In this section, we revisit the CrI3 system (discussed extensively in Section III.1.2) from a different perspective. As previously mentioned, CrI3 exhibits long-range magnetic ordering from few-layer to monolayer.[16] Bulk and few-layer CrI3 possess long-range interlayer and short-range intralayer interactions. The long-range vdW interlayer (noncovalent) interactions, which are due to strongly correlated electrons occupying orbitals of Cr, and the competing intralayer correlations make this an extremely challenging system to model with approaches such as local or semilocal DFT. Certain vdW corrections in DFT can improve the accuracy of these interlayer forces, but the results can vary drastically depending on which vdW correction is employed.
In this work, we studied the binding properties of bulk CrI3 with DMC, with the goal of simultaneously describing the short- and long-range correlations and overcoming the shortcomings of DFT.[342] For the monoclinic bulk CrI3 structure, we calculated the interlayer separation distance to be 6.749(73) Å with DMC, which is in excellent agreement with the experimental value of 6.623 Å.[343] We also estimated the interlayer binding energy to be between 14.3 and 17.9 meV/Å2. We benchmarked several DFT functionals with and without vdW corrections against our DMC results for interlayer separation distance and binding energy (see Fig. 18), and we found that vdW-DF-optB88 [79] and vdW-DF-optB86b [344] are closest at reproducing our DMC values. In addition, we studied the bulk rhombohedral structure of CrI3 and found that the rhombohedral and monoclinic are within thermal energy differences of each other, which is in agreement with experiment.[343] We believe our DMC benchmark calculations can be useful for testing other promising current and future vdW-corrected functionals. Further information about this work can be found in Ref 342.
@@ III.3.6 Graphene-Supported Pt Layers
Recent experimental realization of 2D layered structures of Pt atoms above a graphene surface[345, 346, 347] has opened an interesting avenue for exploration of their various morphologies, as well as the bonding nature of Pt-graphene complexes. In the sense that the layer formation is a consequence of a competition between nondirectional metallic bondings among Pt atoms and their covalent and vdW bondings to graphene, DMC benchmarking is critical for accurate theoretical investigation of these layered systems with large degrees of freedom (including different Pt/C atomic ratios). As far as their geometries and relative energetics between different layered structures of Pt atoms are concerned, DFT results based on the SCAN+rVV10 functional are found to best agree with the corresponding DMC results.
Our DMC-benchmarked DFT calculations show that a (111)-packing structure, where each of the Pt atoms is adsorbed at a hollow site of graphene (see Fig. 19(c)), is more stable than the (100)-packing ones when forming the commensurate Pt-graphene complex at a Pt/C atomic ratio of 1/2. This can be understood by a significant lattice mismatch (>10) between pristine graphene and a freestanding (100)-packing Pt layer. Energetic stability of incommensurate Pt layers with different Pt/C atomic ratios was also investigated. Fig. 19 displays DFT-SCAN+rVV10 formation energies of both flat and buckled incommensurate (100)-packing Pt monolayers (see Fig. 19(a)) as functions of the Pt/C atomic ratio. The lowest-energy configuration is found at the ratio of 0.457, and its formation energy is lower than the corresponding energies for buckled (100)-packing monolayer with the atomic ratio of 1/2, which is in line with recent experimental findings.[347] However, the formation energy of this optimal incommensurate structure is still higher than that of the (111)-packing commensurate structure. Through its systematic investigation for various possible morphologies of layered Pt-on-graphene systems, this DMC-benchmarked study contributes to the expansion of a family of available metallic Pt layers. Additional information can be found in Ref. 348.